Please provide answers for A - C - D (B is correct)Modern Artifacts can produce keepsakes that will be sold for $60 each. Nondepreciation fixed costs are $2,600 per year, and variable costs are $30 per unit. The initial investment of $5,000 will be depreciated straight-line over its useful life of 5 years to a final value of zero, and the discount rate is 10%. a. What is the accounting break-even level of sales if the firm pays no taxes? (Do not round intermediate calculations. Round your answer to the nearest whole number.) Acounting break-even level of sales units b. What is the NPV break-even level of sales if the firm pays no taxes? (Do not round intermediate calculations. Round your answer to the nearest whole number.) NPV break-even level of sales131 units C. what is the accounting break-even level of sales if the firms tax rate is 40%? (Do not round intermediate calculations. Round your answer to the nearest whole number.) Acounting break-even level of sales units d. what is the NPV break-even level of sales if the firms tax rate is 40%? (Do not round intermediate calculations. Round your answer to the nearest whole number.) NPV break-even level of sales units
Answer #1

a.

Contribution per unit = S.P - V.C = 60-30 = 30

Fixed costs = non depreciation costs + depreciation costs = 2600 + 1000 = 3600

Accounting break even = Fixed costs / contribution per unit = 3600/30 = 1,200

c.

Taxes do not change the accounting break-even point, hence the answer is the same 1,200

d.

At break,-even NPV, present value of cash outflows are equal to present value of cash inflows,

As depreciation is not an actual cash outflow/inflow, it will be ignored,

Let the number of units sold be X,

Contribution after tax 30x * 60% = 18x

Non-depreciation fixed costs after tax = 2600 * 60% = 1560

Tax shield on depreciation = 40% * 1000 = 400

Net present value = 18x - 1560 + 400 = 18x -1160

PVAF(10%,5 years) = 3.7907867694085,

(18x - 1160) * 3.7907867694085 - 5000 = 0

68.23416x - 4397.3126 - 5000 = 0

x = 137.72152 or 138